# Experimental validation of a linear momentum and bluff-body model for high-blockage cross-flow turbine arrays

### v1

@@ 3.3 Free Surface Deformation
Figure 10: Time-averaged streamwise water depth profiles across Turbine B for various $\lambda$ at 10 $\beta=36.7\%$, 10 $\beta=45.0\%$, and 10 $\beta=55.0\%$. The farthest upstream measurement is taken at the center of the channel, whereas depth measurements in the vicinity of the array are directly upstream and downstream of the turbine (figure 3). Depth measurements at each streamwise location ($h_{x}$) are normalized by the far-upstream water depth ($h$) measured at the same $\beta$ and $\lambda$. The dashed lines indicate the extent of the turbine’s swept area.
@@
The time-average streamwise free surface profile across Turbine B is shown in figure 10 at the same combinations of $\beta$ and $\lambda$ as for the flow fields in figure 8. For all $\beta$ and $\lambda$, a drop in the free surface is observed across the turbine, consistent with the pressure drop across the array from momentum extraction in an open-channel flow (Vogel et al., 2016). The magnitude of this drop (in this case, the difference between the measured water depths at $x/D=-1$ and $x/D=1$) tends to increase with $\beta$ for a given $\lambda$. However, as shown in Appendix A, the free surface drop behind the array also depends on the normalized submergence depth ($s/h$) which decreases with $\beta$ in these experiments (Section 2.2). Consequently, the free surface drop at higher $\beta$ is likely augmented by the closer proximity of the turbines to the free surface. The free surface drop across the turbine clearly increases with $\lambda$ at $\beta=55.0\%$ (figure 10), consistent with the monotonic increase in $C_{T}$ with $\lambda$ (figure 6). However, even though similar relationships between $C_{T}$ and $\lambda$ are observed for $\beta=36.7\%$ and $\beta=45.0\%$, the measured free surface drop at these blockages does not vary significantly with the tip-speed ratio for $\lambda\geq 2.6$, although the measured depth at $x/D=0.5$ does continuously decrease with the $\lambda$ for all $\beta$.
@@
4 Evaluation of LMADT Accuracy for a Cross-flow Turbine Array
4.1 Model Overview
@@
Figure 11: Linear momentum model for an actuator disk in open-channel flow defined by Houlsby et al. (2008) and adapted to the cross-flow turbine array nomenclature.
@@
The performance, flow field, and free surface measurements described in Section 3 allow us to directly assess the accuracy of a one-dimensional linear momentum actuator disk theory (LMADT) description of the high confinement cross-flow turbine array. Given the equivalence between local and global blockage in the experimental setup (Section 2.3) and appreciable free surface deformation (figure 10), the open-channel LMADT model introduced by Houlsby et al. (2008) is most relevant. We note that, in the limiting case of negligible free surface drop across the turbine array ($Fr_{h}^{2}\rightarrow 0$), this formulation reduces to the well-known model of Garrett & Cummins (2007). Detailed derivations of the LMADT equation set are presented in Houlsby et al. (2008) and Houlsby & Vogel (2017).
@@
A schematic of the open-channel model, as applied to the present cross-flow turbine array, is shown in figure 11. The cross-flow turbine array is represented as a single actuator disk with the same blockage ratio and thrust coefficient as the array. For non-uniform turbine spacing, a theory that represents the individual turbines in the array as separate actuator disks would be required (e.g., Nishino & Willden, 2012b; Vogel et al., 2016; Dehtyriov et al., 2021). The flow upstream of the turbine is unidirectional, uniform, and subcritical with a velocity $U_{\infty}$ and a depth $h$. In response to momentum extraction by the turbine, the “core” flow in the streamtube that passes through the rotor slows from $U_{\infty}$ to $u_{t}$ just upstream of the rotor plane, and further decelerates to $u_{w}$ downstream of the turbine in the near-wake. To satisfy continuity, the fluid outside of this streamtube is accelerated from $U_{\infty}$ to $u_{b}$. The core wake velocity ($u_{w}$) and bypass velocity ($u_{b}$) are defined at a location downstream of the turbine where the pressure in the core and bypass flows is equal and hydrostatic. The flow is assumed to be axisymmetric and inviscid up to this location, downstream of which $u_{w}$ and $u_{b}$ mix. Far downstream, the flow is once again uniform, with reduced depth due to the energy extracted by the disk but higher velocity to maintain constant mass flux through the channel. This model qualitatively captures aspects of the experimentally observed wake and bypass velocities (figure 8). However, although LMADT models are widely used to represent turbine hydrodynamics in confined flows—most often as the basis for analytical blockage corrections (e.g., Bahaj et al., 2007; Whelan et al., 2009; Ross & Polagye, 2020a)—several of the underlying assumptions are violated. For example, the wake is assumed to be axisymmetric and to not mix with the bypass until static pressure has equilibrated between the core and bypass flows.
@@
The implementation of the open-channel model used here follows that of Ross & Polagye (2020a), who reorganized Houlsby et al.’s original formulation into two equations from which $u_{w}$ and $u_{b}$ can be obtained numerically if $U_{\infty}$, $C_{T}$, $Fr_{h}$, and $\beta$ are known:
@@ 6. $\mathbf{x_{5}}$
where $h_{1}$ through $h_{5}$ and $U_{1}$ through $U_{5}$ are the water depth and velocity, respectively, at each station. Once $u_{w}$, $u_{b}$, and $u_{t}$ have been calculated from Eqs. 8, 9 and 10, then $h_{4}$, $h_{3}$, and $h_{2}$ may be calculated following Houlsby et al. (2008) as
@@
In our experiments, because the static pressure equilibrium point is unknown, $x_{4}$ is ambiguous, as are $x_{2}$ and $x_{3}$ due to the non-zero rotor “thickness”. In addition, the flume has insufficient length for the flow to fully mix downstream of the array, such that $x_{5}$ cannot be measured. Consequently, although $u_{b}$, $u_{w}$, and $u_{t}$ may be estimated from available measurements and $h_{2}$ through $h_{5}$ calculated, the exact streamwise positions of $x_{2}$ through $x_{5}$ relative to the experimental array are ambiguous.
4.2 Comparison between Experimental and Modeled Near-Wake Velocities
@@
Figure 12: The 12 bypass velocities, and 12 wake velocities (all normalized by the freestream velocity) predicted by the open-channel LMADT model of Houlsby et al. (2008) from measurements of array thrust and flow conditions.
Using the measured values of $C_{T}$, $U_{\infty}$, $Fr_{h}$, and $\beta$, $u_{b}$ and $u_{w}$ are calculated at each $\lambda$ using Eqs. 8 and 9. The resulting analytically-predicted velocities are normalized by the freestream velocity in figure 12. Qualitatively, trends in near-wake velocities predicted by LMADT agree with trends in the experimental near-wake velocities (figure 8). At a given tip-speed ratio, both $u_{b}$ (figure 12) and $u_{w}$ (figure 12) increase with $\beta$ because confinement accelerates the flow through and around the turbine. As $\lambda$ increases, $u_{b}$ increases at all $\beta$ since thrust increases with $\lambda$ and more flow is diverted around the array, whereas $u_{w}$ decreases with $\lambda$ at all $\beta$ due to momentum extraction by the array. In agreement with experimental trends in the velocity profiles in figure 9, modeled $u_{b}$ increases more strongly with tip-speed ratio at higher blockage. We note that the modeled $u_{w}$ is always positive, indicating that the upper limit of the thrust coefficient for physically-valid solutions to open-channel LMADT (analogous to $C_{T}\leq 1$ for LMADT in unconfined flow) is not exceeded in these experiments.
@@
Figure 14: Time-averaged streamwise velocity profiles measured $0.6D$ downstream of the array normalized by the modeled bypass velocity at 14 $\beta=36.7\%$, 14 $45.0\%$, and 14 $55.0\%$ and various $\lambda$. The profiles shown correspond to the average velocity along the middle $70\%$ of the turbine blade span. Vertical gray rectangles correspond to the bypass region and delineate the relative turbine position and wake flow region (white).
In figure 13, the time-averaged streamwise near-wake velocities at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ and optimal $\lambda$ are normalized by $u_{b}$ and $u_{w}$, respectively. Blue hues correspond to experimental velocities lower than the LMADT-modeled value, red hues correspond to experimental velocities higher than the LMADT-modeled value, and white indicates agreement between the experimental and modeled velocities. The measured bypass velocity is well-predicted by LMADT, as indicated by the substantive white regions in figures 13, 13 and 13. As shown in the bypass-normalized velocity profiles in figure 14, the modeled $u_{b}$ tends to be slightly higher than experimental measurements at the array centerline, but are within approximately $8\%$ of the measured value for all $\beta$ and $\lambda$. In contrast, the measured wake velocities are lower than modeled wake velocities at all $\lambda$, and there is no substantive region of the flow that is well represented by $u_{w}$ (figures 13, 13 and 13). Flow fields normalized by $u_{b}$ and $u_{w}$ at other $\beta-\lambda$ combinations are provided as supplementary material.
@@
Although the time-averaged flow fields in figure 8 show distinct wake and bypass regions downstream of the array, the quantitative agreement between the modeled and measured bypass velocities is remarkable given the multiple violations of LMADT in the near-wake flow field. Examples of these non-idealities are highlighted at optimal tip-speed ratio for $\beta=36.7\%$, $45.0\%$ and $55.0\%$ in figure 15. First, unlike the uniform and one-dimensional flow assumed by the LMADT model, the flow in the near-wake varies in space and time, even within the bypass and wake regions (figures 15, 15 and 15). Second, the flow is also rotational (figures 15, 15 and 15): the periodic hydrodynamics of cross-flow turbines give rise to dynamic stall vortices that are shed into the region between the two turbines, and strong tip vortices form along the top and bottom of the turbine projected area. The asymmetric vorticity produced by these sources is expected to drive flow toward the center of each turbine, contributing to wake asymmetry (Bachant & Wosnik, 2015; Ryan et al., 2016). Third, as suggested by the velocity profiles in figure 9 and the turbulent kinetic energy (TKE) in the near-wake (figures 15, 15 and 15), mixing occurs within the shear layer between the bypass flow and the turbine wake, consistent with a co-flowing turbulent jet. In contrast, the LMADT model assumes that the core and bypass flows do not mix until further downstream from where $u_{b}$ is defined (figure 11). This near-wake mixing necessarily reduces the bypass velocity, which may explain why the modeled bypass velocity is slightly higher than the measured bypass velocity at the array centerline (figure 14). Finally, as described in Section 4.1, $u_{b}$ is specified at a location downstream of the turbines where the static pressure is equal in the wake and bypass flows. Although it is unknown whether this condition is satisfied $0.6D$ downstream of the array, $u_{b}$ is less descriptive of the measured bypass flow further downstream ($1.5D$, provided as supplementary material).
4.3 Comparison between Experimental and Modeled Free Surface Deformation
@@
Figure 16: Modeled free-surface profile at various stations along the channel predicted using the open-channel LMADT model of Houlsby et al. (2008) and the measured $C_{T}$, $h$, and $U_{\infty}$ for 16 $\beta=36.7\%$, 16 $\beta=45.0\%$, and 16 $\beta=55.0\%$ for various $\lambda$. The dashed black line indicates the relative location of the array, represented in the linear momentum model as a thin actuator disk.
For comparison with the measured free surface profiles in figure 10, the theoretical free surface profiles are shown for the same $\beta$ and $\lambda$ in figure 16. As noted in Section 4.1, we emphasize that the exact locations of $x_{2}-x_{4}$ relative to the physical array are ambiguous and that $x_{5}$ is likely further downstream than can be measured given the finite length of the flume downstream of the turbines. Consequently, the locations of $x_{1}$ through $x_{5}$ relative to the actuator disk model of the array shown in figure 16 are qualitative. With these caveats, we consider the measured trends in free surface deformation relative to trends in the modeled values. For all $\beta$, the modeled water depth just upstream of the actuator disk ($h_{2}/h$) increases continuously with $\lambda$. Similarly, the immediate static head drop across the actuator disk ($(h_{3}-h_{2})/h$) and the total free surface drop across the channel ($\Delta h/h$) increase continuously with both $\lambda$ and $\beta$.
This analytical result is consistent with a direct relationship between the pressure drop across the array (i.e., the thrust coefficient) and a drop in the free surface across the array, but conflicts with measured free surface profiles in figure 10. Although, as previously noted, the measured $C_{T}$ increases continuously with $\lambda$ at all $\beta$, the measured free surface drop across the turbines at $\beta=36.7\%$ (Figure 10) and $45.0\%$ (Figure 10) does not depend significantly on $\lambda$ for $\lambda\geq 2.6$. This may suggest that LMADT is not fully descriptive of the evolution of the free surface in the vicinity of high blockage turbines, although the limited spatial and temporal resolution of the instrumentation employed precludes a more comprehensive characterization of the free surface dynamics in this data set. Finally, we do experimentally observe trends consistent with modeled deformation for $\beta=55.0\%$.
4.4 Bluff-Body Scaling of Performance Data

### v2

@@ III.3 Free Surface Deformation
Figure 10: Time-averaged streamwise water depth profiles for various $\lambda$ at 10 $\beta=36.7\%$, 10 $\beta=45.0\%$, and 10 $\beta=55.0\%$. The farthest upstream measurement is taken at the center of the channel, whereas depth measurements in the vicinity of the array are directly upstream and downstream of the Turbine B (i.e., $y/D=0$ in the PIV flow fields). Depth measurements at each streamwise location ($h_{x}$) are normalized by the far-upstream water depth ($h$) measured at the same $\beta$ and $\lambda$. The dashed lines indicate the extent of the turbine’s swept area.
@@
The time-average streamwise free surface profile across Turbine B is shown in Figure 10 at the same $\beta$ and $\lambda$ as for the flow fields in Figure 8. For all $\beta$ and $\lambda$, a free surface drop is observed across the turbine, corresponding to the pressure drop across a turbine in an open-channel flow [6]. The magnitude of this drop (in this case, the difference between the measured water depths at $x/D=-1$ and $x/D=1$) tends to increase with $\beta$ for a given $\lambda$. However, as shown in Appendix A, the free surface drop behind the array also depends on the normalized submergence depth ($s/h$) which decreases with $\beta$ in these experiments (Section II.2). Consequently, the free surface drop at higher $\beta$ is likely augmented by the closer proximity of the turbines to the free surface. The free surface drop across the turbine clearly increases with $\lambda$ at $\beta=55.0\%$ (Figure 10), consistent with the monotonic increase in $C_{T}$ with $\lambda$ (Figure 6). However, even though similar relationships between $C_{T}$ and $\lambda$ are observed for $\beta=36.7\%$ and $\beta=45.0\%$, the free surface drop at these blockages does not vary significantly with the tip-speed ratio for $\lambda\geq 2.6$, although the depth at $x/D=0.5$ does continuously decrease with $\lambda$ for all $\beta$.
@@
IV Evaluation of LMADT for a Cross-flow Turbine Array at High Blockage
IV.1 Model Overview
@@
Figure 11: Linear momentum model for an actuator disk in open-channel flow defined by Houlsby et al. [1] and adapted to the cross-flow turbine array nomenclature.
@@
The performance, flow field, and free surface measurements described in Section III allow us to directly assess the accuracy of a one-dimensional linear momentum actuator disk theory (LMADT) description of the cross-flow turbine array. Given the appreciable free surface deformation (Figure 10) and the homogeneous lateral distribution of turbines in the channel (i.e., equivalent local and global blockage), the open-channel LMADT model introduced by Houlsby et al. [1] is most relevant. In the limiting case of negligible free surface drop across the turbine array ($Fr_{h}^{2}\rightarrow 0$), Houlsby et al.’s model reduces to the well-known model of Garrett and Cummins [4]. For a non-uniform distribution of turbines along the channel width, a theory that represents the individual turbines in the array as separate actuator disks would be required [e.g., 5, 6, 8]. Although the array’s homogeneity does not extend to the vertical direction, the blockage asymmetry in these experiments (1.3 - 2.3) is within the range found by Kinsey and Dumas [17] to have limited influence on turbine power and thrust. We note that LMADT-based models like that of Houlsby et al. [1] are only one of several types of analytical models (e.g., double-multiple streamtube models [62, 63], potential flow models [24]) that could be applied. The potential flow model of Steiros et al. [24] is given separate consideration in supplementary material [59].
@@
In the LMADT model, the array is represented as a single actuator disk with the same blockage ratio and thrust coefficient as the array (Figure 11). The flow upstream of the turbine is unidirectional, uniform, and subcritical (i.e., $Fr_{h}^{2}<1$) with velocity $U_{\infty}$ and depth $h$. In response to momentum extraction by the turbine, the “core” flow in the streamtube that passes through the rotor decelerates from $U_{\infty}$ to $u_{t}$ just upstream of the rotor plane, and further decelerates to $u_{w}$ downstream of the turbine in the near-wake. To satisfy continuity, the fluid outside of this streamtube is accelerated from $U_{\infty}$ to $u_{b}$. The core wake velocity ($u_{w}$) and bypass velocity ($u_{b}$) are both defined at a location downstream of the turbine where the pressure in the core and bypass flows is equal and hydrostatic. The flow is assumed to be axisymmetric and inviscid up to this location, downstream of which $u_{w}$ and $u_{b}$ mix. Far downstream, the flow is once again uniform, with reduced depth due to the energy extracted by the disk but higher velocity to conserve mass. This model qualitatively captures aspects of the experimentally observed wake and bypass velocities (Figure 8). However, although LMADT models are widely used to represent turbine hydrodynamics in confined flows, several of the underlying assumptions are often violated. For example, the wake is assumed to be axisymmetric and to not mix with the bypass until static pressure has equilibrated between the core and bypass flows, while we observe a non-uniform wake with immediate mixing between core and bypass.
@@
Detailed derivations of the LMADT equation set used here are presented in Houlsby et al. [1] and Houlsby and Vogel [64]. The implementation in this study follows Ross and Polagye [23], who reorganized Houlsby et al.’s original formulation into two equations from which $u_{w}$ and $u_{b}$ can be obtained numerically if $U_{\infty}$, $C_{T}$, $Fr_{h}$, and $\beta$ are known:
@@ 6. $\mathbf{x_{5}}$
where $h_{1}$ through $h_{5}$ and $U_{1}$ through $U_{5}$ are the water depth and velocity, respectively, at each station. Once $u_{w}$, $u_{b}$, and $u_{t}$ have been calculated from Eqs. 8, 9 and 10, then $h_{4}$, $h_{3}$, and $h_{2}$ may be calculated following Houlsby et al. [1] as
@@
In our experiments, because the static pressure equilibrium point is unknown, the location of $x_{4}$ is ambiguous, as are the locations of $x_{2}$ and $x_{3}$ due to the non-zero rotor “thickness”. In addition, the flume has insufficient length for the flow to fully mix downstream of the array, such that $x_{5}$ cannot be measured. Consequently, although $u_{b}$, $u_{w}$, and $u_{t}$ may be estimated from available measurements and $h_{2}$ through $h_{5}$ calculated, the corresponding streamwise positions of $x_{2}$ through $x_{5}$ in experiments are uncertain.
IV.2 Comparison between Experimental and Modeled Near-Wake Velocities
@@
Figure 12: The 12 bypass velocities, and 12 wake velocities (all normalized by the freestream velocity) predicted by the open-channel LMADT model of Houlsby et al. [1] from measurements of array thrust and flow conditions.
Using the measured values of $C_{T}$, $U_{\infty}$, $Fr_{h}$, and $\beta$, $u_{b}$ and $u_{w}$ are calculated at each $\lambda$ using Eqs. 8 and 9. The resulting analytically-predicted velocities are normalized by the freestream velocity in Figure 12. At a given tip-speed ratio, both $u_{b}$ (Figure 12) and $u_{w}$ (Figure 12) increase with $\beta$ because confinement accelerates the flow through and around the turbine. $u_{b}$ increases with $\lambda$ at all $\beta$ since the corresponding increase in thrust diverts more flow around the array, whereas $u_{w}$ decreases with $\lambda$ at all $\beta$ due to momentum extraction by the array. In agreement with experimental trends in the velocity profiles in Figure 9, modeled $u_{b}$ increases more strongly with tip-speed ratio at higher blockage. We note that the modeled $u_{w}$ is always positive, indicating that the upper limit of the thrust coefficient for physically-valid solutions to open-channel LMADT (analogous to $C_{T}\leq 1$ for LMADT in unconfined flow) is not exceeded in these experiments.
@@
Figure 14: Time-averaged streamwise velocity profiles measured $0.6D$ downstream of the array normalized by the modeled bypass velocity at 14 $\beta=36.7\%$, 14 $45.0\%$, and 14 $55.0\%$ and various $\lambda$. The profiles shown correspond to the average velocity along the middle $70\%$ of the turbine blade span. Vertical gray rectangles correspond to the bypass region.
In Figure 13, the time-averaged near-wake velocity fields at $\beta=36.7\%$, $45.0\%$, and $55.0\%$ and optimal $\lambda$ are normalized by $u_{b}$ and $u_{w}$, respectively. Blue hues correspond to experimental velocities lower than the LMADT-modeled value, red hues correspond to experimental velocities higher than the LMADT-modeled value, and white indicates agreement between the experimental and modeled velocities. The measured bypass velocity is well-predicted by LMADT, as indicated by the substantive white regions in Figures 13, 13 and 13. As shown in the bypass-normalized velocity profiles in Figure 14, the modeled $u_{b}$ tends to be slightly higher than experimental measurements at the array centerline, but are within approximately $8\%$ of the measured value for all $\beta$ and $\lambda$. In contrast, the measured wake velocities are lower than modeled wake velocities at all $\lambda$, and there is no substantive region of the flow that is well represented by $u_{w}$ (Figures 13, 13 and 13). Flow fields normalized by $u_{b}$ and $u_{w}$ at other $\beta-\lambda$ combinations are provided as supplementary material [59]. While it is unknown if the static pressure is equal in the wake and bypass flows at $0.6D$ downstream of the array, $u_{b}$ is less descriptive of the measured bypass flow further downstream ($1.5D$), as shown in supplementary material [59]).
@@
The quantitative agreement between the modeled and measured bypass velocities is remarkable given the multiple violations of LMADT in the near-wake flow field. First, unlike the uniform and one-dimensional flow assumed by the LMADT model, the flow in the near-wake varies in space and time, even within the bypass and wake regions (Figures 15, 15 and 15). Second, the flow is rotational (Figures 15, 15 and 15): dynamic stall vortices are shed into the region between the two turbines and strong tip vortices form along the top and bottom of the turbine projected area. The asymmetric vorticity produced by these sources drives flow toward the center of each turbine, contributing to wake asymmetry [65, 66]. Third, as suggested by the velocity profiles in Figure 9 and the turbulent kinetic energy (TKE) in the near-wake (Figures 15, 15 and 15), mixing occurs between the bypass flow and the turbine wake, consistent with the shear layer in a co-flowing turbulent jet. In contrast, the LMADT model assumes that the core and bypass flows do not mix until further downstream (Figure 11). This near-wake mixing necessarily reduces the bypass velocity, which may explain why the modeled bypass velocity is slightly higher than the measured bypass velocity at the array centerline (Figure 14).
IV.3 Comparison between Experimental and Modeled Free Surface Deformation
@@
Figure 16: Modeled free-surface profile at various stations along the channel predicted using the open-channel LMADT model of Houlsby et al. [1] and the measured $C_{T}$, $h$, and $U_{\infty}$ for 16 $\beta=36.7\%$, 16 $\beta=45.0\%$, and 16 $\beta=55.0\%$ for various $\lambda$. The dashed black line indicates the relative location of the array, represented in the linear momentum model as a thin actuator disk.
For comparison with the measured free surface profiles in Figure 10, the theoretical free surface profiles are shown for the same $\beta$ and $\lambda$ in Figure 16. As noted in Section IV.1, we emphasize that the exact locations of $x_{2}-x_{4}$ relative to the physical array are ambiguous and that $x_{5}$ is likely further downstream than can be measured given the finite flume length. Consequently, the locations of $x_{1}$ through $x_{5}$ relative to the actuator disk model of the array shown in Figure 16 are qualitative. With these caveats, we consider the trends in measured free surface deformation relative to model predictions. For all $\beta$, the modeled water depth just upstream of the actuator disk ($h_{2}/h$) increases continuously with $\lambda$. Similarly, the immediate static head drop across the actuator disk ($(h_{3}-h_{2})/h$) and the total free surface drop along the channel ($\Delta h/h$) increase continuously with both $\lambda$ and $\beta$.
This analytical result is consistent with a direct relationship between the pressure drop across the array (i.e., the thrust coefficient) and a drop in the free surface across the array, but conflicts with measured free surface profiles in Figure 10. Although the measured $C_{T}$ increases continuously with $\lambda$ at all $\beta$, the measured free surface drop across the turbines at $\beta=36.7\%$ (Figure 10) and $45.0\%$ (Figure 10) does not depend significantly on $\lambda$ for $\lambda\geq 2.6$. This may suggest that LMADT is not fully descriptive of the evolution of the free surface in the vicinity of high blockage turbines, although the limited spatial and temporal resolution of the instrumentation employed precludes a more comprehensive characterization of the free surface dynamics in this data set. Experimental trends are consistent with modeled free surface deformation for $\beta=55.0\%$, which suggests that smaller-scale spatial variability could be masking expected trends at lower $\beta$.
IV.4 Bluff-Body Scaling of Performance Data
Given that open-channel LMADT is descriptive of the measured bypass flow across a range of array operating conditions, we turn to the question of whether LMADT also describes the measured array performance. In prior work, LMADT is commonly used in analytical blockage corrections to estimate unconfined performance from measurements or simulations of confined operation [e.g., 67, 20, 17, 23]. However, since a perfect blockage correction would be expected to collapse the $C_{P}$ and $C_{T}$ measured at different blockage ratios (e.g., Figure 6) onto common unconfined $C_{P}-\lambda$ and $C_{T}-\lambda$ curves, such corrections imply not just a connection between confined and unconfined performance, but rather a broader self-similarity in turbine performance across blockage ratios. Here, we explore such self-similar aspects of array performance with a focus on the significance of the bypass flow.